OpenStudy (anonymous):

solve: 2sqrt. x =4 sqrt. 3

OpenStudy (anonymous):

\[2\sqrt{x}=4\sqrt{3}\]

OpenStudy (anonymous):

Divide both sides by 2. \[\sqrt{x}=2\sqrt{3}\] Square both sides: \[x=4(3)\] \[x=12\]

OpenStudy (anonymous):

To prove this: \[2\sqrt{12}=2\sqrt{4*3}=4\sqrt{3}\]

OpenStudy (anonymous):

:D thank you very much

OpenStudy (anonymous):

Welcome.